on a door is a sign that says WELCOME TO OUR HOME.the sign is in the shape of a regular octogon.what is the measure of one of the interior angeles of the sign?
Answer by user_dude2008(1862)

Use formula: (180(n-2))/n .... n = number of sides

Interior Angle = (180(n-2))/n = (180(8-2))/8 = (180(6))/8 = 1080/8 = 135


Ans: Interior Angle = 135 degrees